The sum of a number and its square is 30. Find the​ number(s).
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The sum of a number and its square is 30. Find the​ number(s).
-----
Let the number be "x"::
Equation:
x + x^2 = 30
x^2 + x - 30 = 0
---
Factor::
(x-6)(x+5) = 0
----
x = 6 or x = -5
------------------
Cheers,
Stan H.
--------

Answer by Alan3354(69443)

x^2 + x = 30
z^2 + x - 30 = 0
(x-5)*(x+6) = 0
x = -6, x = 5